(a) In 2000, the scores of students taking SATs were normally distributed with mean 1019 and standard deviation 209. (i) What percent of all students had the SAT scores of at least 820? [3 marks] (ii) What percent of all students had the SAT scores between 720 and 820? [3 marks] (iii) How high must a student score in order to place in the top 20% of all students taking the SAT? [4 marks]
Answer #1

\mu =1019,\sigma =209

i) Z=\frac{X-\mu }{\sigma }=\frac{820-1019}{209}=\frac{-199}{209}=-0.95

P(X\geq 820)=P(Z\geq -0.95)=0.8295

ii)Z_{1}=\frac{X_{1}-\mu }{\sigma }=\frac{720-1019}{209}=\frac{-299}{209}=-1.43

Z_{2}=\frac{X_{2}-\mu }{\sigma }=\frac{820-1019}{209}=\frac{-199}{209}=-0.95

P(720< X< 820)=P(-1.43< Z< -0.95)=0.0942

iii) Z-score of top 20% is Z=0.84

Let required score be X

X=\mu +Z\sigma =1019+0.84\times 209=1195